OpenStudy (anonymous):

PLEASE HELP ANYONE!!! The number of bacteria p in a pond after t hours is p(t). The more bacteria there are, the more rapidly the number increases. This is expressed by the equation p'(t)=6p(t). Find the most general exponential function of the form A times e^Kt which satisfies the equation. (A and K are constants. Hint: substitute this function into the equation and see what you get). p(t)= If there were 1000 bacteria at t=0, how many are there 3 hours later? p(3)=

zepdrix (zepdrix):

Hmm so this is a separable differential equation. This will work out maybe a little different than you're used to. :o

zepdrix (zepdrix):

\[\Large \frac{dp}{dt}\quad=\quad 6p\]

zepdrix (zepdrix):

Let's ignore the hint for now and just do it the legit way XD

zepdrix (zepdrix):

So we want to start by `separating` the variables. We want to get the `p's and dp` on one side, and the `t's and dt` on the other. So we'll "multiply" the dt to the other side. And divide both sides by p.\[\Large \frac{dp}{p}=6dt\]

zepdrix (zepdrix):

See how we have dp and p on the same side now? :O Making a little bit of sense maybe? :x

zepdrix (zepdrix):

So our next step is to integrate.\[\Large \int\limits\frac{dp}{p}=\int\limits6dt\]

zepdrix (zepdrix):

Where you at sam? XD wake up!!

OpenStudy (anonymous):

im here lol

OpenStudy (anonymous):

are you still there?

zepdrix (zepdrix):

So what do we get when we integrate the p side? :D

zepdrix (zepdrix):

\[\Large \int\limits \frac{1}{p}dp\quad=\quad?\]

OpenStudy (anonymous):

ummm.. 1

zepdrix (zepdrix):

No we can't apply the power rule to a term with a -1 exponent. Can you think of a special function that gives us 1/x when we take it's derivative?

OpenStudy (anonymous):

its not 1

OpenStudy (anonymous):

im lost honestly

zepdrix (zepdrix):

\[\Large (\ln x)' \quad=\quad \frac{1}{x}\]Right? +_+

OpenStudy (anonymous):

yes yes i see

zepdrix (zepdrix):

So for the left side,\[\Large \color{royalblue}{\int\limits\limits\frac{dp}{p}}=\int\limits\limits6dt\]We'll get,\[\Large \color{royalblue}{\ln p+C}=\int\limits\limits\limits6dt\]

zepdrix (zepdrix):

We add a +C when we do indefinite integrals :D

zepdrix (zepdrix):

How bout the integral on the right side?\[\Large 6\int\limits \;dt \quad=\quad ?\]

OpenStudy (anonymous):

ummm...

OpenStudy (anonymous):

im sorry my brain isnt as sharp as it should be since its like 11:16 here

zepdrix (zepdrix):

XD fair

zepdrix (zepdrix):

\[\Large \ln p+C=6\color{royalblue}{\int\limits(1)dt}\]When we integrate 1, with respect to t, it becomes t. You can always take the derivative to check. Derivative of t is 1, yes? :x \[\Large \ln p+C=6\color{royalblue}{(t+c)}\]

OpenStudy (anonymous):

yes

zepdrix (zepdrix):

Distribute the 6,\[\Large \ln p+C=6t+c\](Absorb the 6 into the c, cuz whatev, it doesn't matter, c is arbitrary).

zepdrix (zepdrix):

Subtract C from each side,\[\Large \ln p\quad=\quad 6t+\color{orangered}{c-C}\]Again, since the constants are arbitrary (could take on any value), we can combine them into a new constant.\[\Large \ln p\quad=\quad 6t+\color{orangered}{d}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

We want to solve for p. So from here we'll have to do a sneaky little math trick. We'll `exponentiate` each side. Rewrite each side as an exponent with base e.\[\LARGE e^{\ln p}\quad=\quad e^{6t+d}\]

zepdrix (zepdrix):

Since the `exponential function` and the `log function` are inverses of one another, they "undo" each other. So we simply get p on the left side.\[\LARGE p=e^{6t+d}\]

zepdrix (zepdrix):

From here we want to apply a rule of exponents:\[\Large \color{teal}{e^{a+b}\quad=\quad e^a\cdot e^b}\]

zepdrix (zepdrix):

So applying this rule gives us:\[\LARGE p\quad=\quad e^{6t}\cdot e^d\]

zepdrix (zepdrix):

Again, we're going to do something annoying :) e^d is simple a weird looking constant, we'll label it as something a little more manageable\[\Large p\quad=\quad A\cdot e^{6t}\].

zepdrix (zepdrix):

Soooo I think that's the form they were looking for +_+ That takes care of the first part.

OpenStudy (anonymous):

let me see

zepdrix (zepdrix):

\[\Large p(t)\quad=\quad A\cdot e^{6t}\]

OpenStudy (anonymous):

is their possibly another way webwork can accept this answer like the issue earlier?

zepdrix (zepdrix):

Oh hmm I remember that... A*exp(6t) yah try that.

OpenStudy (anonymous):

correct!

zepdrix (zepdrix):

Next part is kinda tricky... we have to use the first piece of information to find A. Then we plug in our A value and find p(t) at t=3.

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

1000 bacteria at t=0.\[\Large p(0)\quad=\quad 1000\qquad\to\qquad Ae^{6(0)}\quad=\quad1000\]

zepdrix (zepdrix):

So what do we get for A?

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

1000^18

zepdrix (zepdrix):

\[\Large Ae^{0}=1000\] \[\Large e^0\quad=\quad?\]

OpenStudy (anonymous):

1

zepdrix (zepdrix):

\[\Large Ae^{0}=1000\]\[\Large A(1)=1000\]\[\Large A=1000\]

OpenStudy (anonymous):

my bad

zepdrix (zepdrix):

\[\Large p(t)\quad=\quad A\cdot e^{6t}\]\[\Large p(t)\quad=\quad1000\cdot e^{6t}\]

OpenStudy (anonymous):

ok

zepdrix (zepdrix):

\[\Large p(3)\quad=\quad1000\cdot e^{6(3)}\]

OpenStudy (anonymous):

1000*e^18

zepdrix (zepdrix):

yes. Your webworks might want you to put it in as a decimal value. Or maybe like this: 1000*exp(18) i'm not sure which.

OpenStudy (anonymous):

it works thanks!!!!!!!!